OpenStudy (bbb911):

Jaci averaged 84 on her first three history exams. The first score was 67. The second score was 7 less than the third score. What did she score on the second and third exams?

jimthompson5910 (jim_thompson5910):

let x = score on second exam and y = score on third exam

OpenStudy (bbb911):

so x would be 67

jimthompson5910 (jim_thompson5910):

we know that "The second score was 7 less than the third score", so we can make this equation: x = y - 7

jimthompson5910 (jim_thompson5910):

67 is the first exam score

jimthompson5910 (jim_thompson5910):

the second may be 67, but we haven't gotten there yet

OpenStudy (bbb911):

so would it be like this 67x=y-7

jimthompson5910 (jim_thompson5910):

so we know this so far first exam score: 67 second exam score: x third exam score: y replace x with y-7 (since x = y-7) to get first exam score: 67 second exam score: y-7 third exam score: y

jimthompson5910 (jim_thompson5910):

now you add up all 3 scores, then divide by 3 to get the average of 84 like this (Score1 + Score2 + Score3)/3 = Average (67 + (y-7) + y)/3 = 84 (67 + y-7 + y)/3 = 84 (2y+60)/3 = 84 2y+60 = 84*3 ... ... ... y = ???

OpenStudy (bbb911):

oh ok so i use this equation right ?:)

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (bbb911):

kk one sec

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (bbb911):

so you don't have to divide the 3?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (bbb911):

(2y+60)/3 = 84 2y+60 = 84*3 like here how instead of dividing you multiplied 3 to 84

jimthompson5910 (jim_thompson5910):

oh i multiplied both sides by 3 because in the end you want to isolate y

jimthompson5910 (jim_thompson5910):

so you undo that division by 3

OpenStudy (bbb911):

oh ok :) thnks one moment

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (bbb911):

is it 96 ?

jimthompson5910 (jim_thompson5910):

yep, y = 96 now find x

OpenStudy (bbb911):

okay now do you plug in the y to the very first equation you gave me?

jimthompson5910 (jim_thompson5910):

yep x = y - 7

OpenStudy (bbb911):

is it 89

jimthompson5910 (jim_thompson5910):

yep x = 89 y = 96

OpenStudy (bbb911):

yesss! thank you so mcuh :)

jimthompson5910 (jim_thompson5910):

so going back to what we defined x and y to be (remember at the top we said "let x = score on second exam and y = score on third exam", so you can see how it all comes around full circle) we know first exam score = 67 (given) second exam score = 89 (just found) third exam score = 96 (just found)

jimthompson5910 (jim_thompson5910):

yw